# Multi Source COVID-19 Detection via Kernel-Density-based Slice Sampling

### v1

@@ 1 Introduction
The COVID-19 pandemic has increased interest in automated diagnostic tools to support clinical decision-making. Chest computed tomography (CT) can reveal pulmonary manifestations such as ground-glass opacities and bilateral infiltrates associated with COVID-19 [6, 7].
Developing automated COVID-19 detection systems from CT scans faces challenges in multi-institutional settings. CT data from different hospitals exhibit variability due to differences in acquisition parameters, scanner models, imaging protocols, and patient demographics. This domain shift can affect deep learning model performance when deployed across different medical centers [9, 8, 10].
To address these challenges, we participate in the Multi-Source COVID-19 Detection Challenge, which requires binary classification of COVID/Non-COVID cases using CT scans from four diverse medical institutions. The heterogeneous nature of this dataset presents unique opportunities to develop and validate domain-robust algorithms that can generalize across different clinical environments.
Our approach builds upon the SSFL framework [4, 3, 2, 5], incorporating Kernel-Density-based Slice Sampling (KDS) to handle multi-source variability effectively. The key insight is that while CT scans from different institutions may vary in acquisition parameters and quality, the underlying anatomical structures and pathological patterns remain consistent. By focusing on diagnostically relevant lung regions and intelligently sampling representative slices, our method can learn domain-invariant features while maintaining computational efficiency.
Through experiments comparing EfficientNet [12] and Swin Transformer [11] architectures, we evaluate the effectiveness of our preprocessing pipeline on multi-source data. Our results contribute to understanding the trade-offs between model complexity and preprocessing design in multi-institutional medical imaging tasks.

### v2

@@ 1 Introduction
Deep learning has become a cornerstone of modern medical–image analysis, demonstrating strong predictive power across diverse healthcare applications [9, 10]. During the COVID-19 pandemic, this momentum spurred rapid development of computer-aided diagnostic systems. Chest computed-tomography (CT) imaging, in particular, can reveal hallmark pulmonary manifestations such as ground-glass opacities and bilateral infiltrates that correlate with infection severity [12, 13]. Early 3-D CNN pipelines trained on large, curated CT datasets confirmed that automated screening could approach radiologist-level sensitivity for COVID-19 detection and severity grading [1].
Despite these successes, multi-institutional deployment remains difficult. CT volumes collected at different hospitals vary in scanner models, reconstruction kernels, slice spacing and patient demographics, introducing a domain shift that can markedly degrade performance when a network is evaluated outside its source domain [15]. Recent studies emphasise that robust COVID-19 screening must explicitly address site heterogeneity through domain adaptation, fairness and explainability mechanisms [14, 16]. Transparent adaptation strategies—which align latent representations while exposing feature changes to clinicians—have been shown to mitigate such shifts without sacrificing interpretability [11].
To probe these issues, we participate in the PHAROS-AFE-AIMI Multi-Source Covid-19 Detection, which provides chest CT scans from four distinct medical centres and requires binary classification of covid versus non-covid cases. The benchmark’s heterogeneous acquisition protocols mirror real-world deployment scenarios and thus constitute a demanding test of domain generalisation [2, 4].
Our approach extends the SSFL framework of Hsu et al.[6] by integrating a Kernel-Density–based Slice-Sampling (KDS) [7] scheme that selects diagnostically relevant slices while preserving anatomical coverage. Because COVID-19 lesions are confined to lung parenchyma regardless of scanner vendor, we first extract precise lung masks with the recent vision–language segmentation pipeline Sam2Clip2Sam [17]. KDS then draws balanced slice subsets conditioned on kernel-density estimates of regional intensity distributions, yielding a compact yet domain-invariant representation for downstream classification. We compare a lightweight CNN (EfficientNet) and a transformer backbone (Swin Transformer) under identical preprocessing, demonstrating that careful slice selection narrows the gap between model sizes while improving cross-site robustness.